# Local asymptotics of selection models with applications in Bayesian selective inference

### v2

@@ Abstract
Contemporary focus on selective inference provokes interest in the asymptotic properties of selection models, as the working inferential models in the conditional approach to inference after selection. In this paper, we derive an asymptotic expansion of the local likelihood ratios of selection models. We show that under mild regularity conditions, they behave asymptotically like a sequence of Gaussian selection models. This generalizes the Local Asymptotic Normality framework of Le Cam (1960) to a class of non-regular models, and indicates a notion of local asymptotic selective normality as the appropriate simplifying theoretical framework for analysis of selective inference. Furthermore, we establish practical consequences for Bayesian selective inference. Specifically, we derive the asymptotic shape of Bayesian posterior distributions constructed from selection models, and show that they will typically be significantly miscalibrated in a frequentist sense, demonstrating that the familiar asymptotic equivalence between Bayesian and frequentist approaches does not hold under selection.
Keywords: Asymptotics; Local asymptotic normality; Frequentist calibration; Posterior distribution; Selective inference; Selection models.
@@ 1 Introduction
The techniques of selective inference aim to ensure validity in situations where the same data is used to select the inference to be considered and also to conduct it. Three main approaches have been proposed: the simultaneous approach (Berk et al., 2013), which ensures validity irrespective of the inference performed and the selection procedure; the ‘condition on selection’ approach (Fithian et al., 2017), where inference is conditioned on those hypothetical datasets which might have occurred which would have led to selection of the same inferential problem; and information splitting methods, such as data splitting, where information in the sample is split between that used for selection and that used for inference: see for example García Rasines and Young (2023).
@@
In this work, we study the asymptotic behavior of selection models, which arise as the working inferential models under the conditional and information-splitting approaches to selective inference. These are statistical models of substantial contemporary practical importance in which the probability of observing or analyzing a sample depends on the sample itself. The term ‘selection model’ was coined by Fraser (1952, 1966), although their study dates back at least to Fisher (1934). Selection models are ubiquitous in statistical inference and have traditionally been used to model situations involving sampling bias, where samples are observed only if they satisfy a certain condition—for example, if they belong to a certain portion of the population (Rao, 1985). Selective inference uses a rather more general formulation, which is described in Section 2.
@@
A significant body of methodological research in selective inference has been developed for Gaussian models, partly because they allow for tractable analytical procedures in some common contexts; see e.g. Lee et al. (2016). As a result, the theoretical properties of selection models within this setting are well understood, both from the frequentist and the Bayesian positions. The main contribution of this work is the identification of an asymptotic connection between general parametric selection models and Gaussian ones, thereby providing a uniform framework for analyzing the properties of the former and therefore deriving powerful theoretical insights.
@@
More broadly, this paper contributes to the literature on non-standard asymptotics by establishing a limiting behavior which deviates from the standard Gaussian framework. We show that, under mild regularity conditions, a sequence of selection models from a parametric distribution is asymptotically equivalent to a sequence of selection models derived from a Gaussian location model, which, in contrast to standard asymptotics, can be highly asymmetrical if the selection effect is significant (Section 3). The relevant Gaussian models inherit a selection mechanism for the mean parameter that often depends on both the original model and the underlying data-generating parameter in an intricate manner, as illustrated in the examples of Sections 3.1 and 3.2. This showcases the unconventional nature of the analysis, compared to standard asymptotic theory, and underscores the need for careful consideration of the appropriate asymptotic framework for inference.
@@
As a natural corollary of the asymptotic approximation, we derive an extension of the Bernstein-von Mises Theorem to selection models (Section 4). We show that in a selection model with a fixed prior on the parameter, the Bayesian posterior is asymptotically equivalent to that obtained in a Gaussian selection model with a uniform prior on the location parameter. Crucially, this is a non-Gaussian limit, with significant practical consequences for selective inference, as the posterior in such a Gaussian selection model is miscalibrated from a frequentist perspective. In these models, frequentist calibration requires specification of a prior density which depends strongly on the sample size, as discussed in García Rasines and Young (2022).
2 Selection models
@@
Let $Y^{n}=[Y_{1},\ldots,Y_{n}]$ be independent and identically distributed (IID) random variables from a distribution $F_{\theta}$ on a measurable space $(\mathcal{Y},\mathcal{A})$, where $\theta$ is a parameter from an open set $\Theta\subseteq\mathbb{R}^{p}$. The statistical model for the full sample, defined over $(\mathcal{Y}^{n},\mathcal{A}^{n})$, is $\mathcal{M}_{n}=\{F_{\theta}^{n}\colon\theta\in\Theta\}$. It is well known that, under mild regularity conditions, $\mathcal{M}_{n}$ can be asymptotically approximated by a Gaussian location model. Specifically, if $\{F_{\theta}\colon\theta\in\Theta\}$ is differentiable in quadratic mean (DQM) (van der Vaart, 1998, p. 93), then as $n\to\infty$ the sequence of local log-likelihood ratios admits the asymptotic expansion
@@
$\log\prod_{i=1}^{n}\frac{f_{\theta+h_{n}/\sqrt{n}}(Y_{i})}{f_{\theta}(Y_{i})}=h^{T}I_{\theta}Z_{n}-\frac{1}{2}h^{T}I_{\theta}h+o_{F_{\theta}}(1)$
@@
for every converging sequence $h_{n}\to h$, where $Z_{n}\rightsquigarrow N(0,I_{\theta}^{-1})$, $I_{\theta}$ is the Fisher information for a single sample, and $f_{\theta}(y)$ is the density of $F_{\theta}$ with respect to some fixed measure $\mu$ (we use $\rightsquigarrow$ to denote convergence in distribution). This property, formalized by Le Cam (1960), is known as Local Asymptotic Normality (LAN). Many key asymptotic results in parametric statistics can be derived from the LAN property, including the asymptotic distribution of maximum likelihood estimators and the Bernstein–von Mises theorem in Bayesian inference; see van der Vaart (1998) for a detailed account.
@@
In wide generality, a selection model can be obtained from a base model $\mathcal{M}_{n}$ by conditioning the data on a random quantity $U_{n}$ such that $U_{n}\mid Y^{n}$ is distribution-constant. Typically, $U_{n}$ is either a function of $Y^{n}$, such as $\mathbf{1}(Y_{1}+\ldots+Y_{n}>0)$, where $\mathbf{1}(A)$ denotes the indicator function of the event $A$, or a randomized function of it, $u(Y^{n},W_{n})$, say, where $W_{n}$ is user-generated noise independent of $\theta$. For a realized value $U_{n}=u_{n}$, the conditional density of $Y^{n}\mid u_{n}$ is
@@
where $f(u_{n}\mid y^{n})$ and $f_{\theta}(u_{n})$ are the conditional and marginal densities of $U_{n}$ relative to some dominating measures. We denote the corresponding probability distribution by $F_{\theta}^{u_{n}}$. In our analysis, we will treat the sequence $\{u_{n}\colon n\in\mathbb{N}\}$ as fixed, as it represents the quantity being conditioned on in the analysis, so we will occasionally suppress $u_{n}$ from the notation. In particular, we will write
@@
where $p_{n}(y^{n})=f(u_{n}\mid y^{n})$ and $\varphi_{n}(\theta)=f_{\theta}(u_{n})$. This simplifies the expressions, in particular in Section 4 where we discuss Bayesian selective inference, and aligns better with the notation conventionally employed in most practical applications of selective inference.
Throughout this paper, the term selection model refers to any statistical model of the form $\{F_{\theta}^{u_{n}}:\theta\in\Theta\}$, derived from an underlying base model $\{F_{\theta}^{n}:\theta\in\Theta\}$ and a specified conditioning statistic $U_{n}$. Furthermore, by a Gaussian selection model is meant a selection model where the base model $F_{\theta}$ is Gaussian with a known variance matrix and an unknown mean $\theta\in\mathbb{R}^{p}$. In a generic selection model, we refer to $p_{n}(y^{n})$ as the selection function This encodes all relevant distributional information about the selection mechanism, in the sense that any two statistics and realized values $U_{n1}=u_{n1}$ and $U_{n2}=u_{n2}$ with proportional selection functions $p_{n1}(y^{n})\propto p_{n2}(y^{n})$ induce the same selection model on $Y^{n}$. We denote the corresponding sequence of selection models by $\mathcal{M}_{n}(u_{n})$.

### v3

@@ Abstract
Contemporary focus on selective inference has renewed interest in the theory of selection models. In this paper, we analyze the asymptotic properties of selection models built on independent and identically distributed observations. We show that, under suitable regularity conditions, they behave asymptotically like a sequence of Gaussian selection models. This provides a natural generalization of the Local Asymptotic Normality framework of Le Cam (1960), and indicates a notion of local asymptotic selective normality as the appropriate simplifying theoretical framework for analysis of selective inference. As a key application, we consider the methodological consequences of the asymptotic theory for Bayesian selective inference. Specifically, we prove that the posterior distribution constructed from a selection model under a fixed prior is asymptotically equivalent to the posterior derived in the corresponding asymptotic Gaussian selection model under a uniform prior. Notably, the latter is often mis-calibrated in a frequentist sense, particularly for one-sided selection mechanisms. This demonstrates that the familiar asymptotic equivalence between Bayesian and frequentist approaches does not hold under selection.
Keywords: Asymptotics, Local asymptotic normality, Frequentist calibration, Posterior distribution, Selective inference, Selection models.
@@ 1 Introduction
The techniques of selective inference aim to ensure validity in situations where the same data is used to select the inference to be considered and also to conduct it. Three main approaches have been proposed: the simultaneous approach (Berk et al., 2013), which ensures validity irrespective of the inference performed and the selection procedure; the ‘condition on selection’ approach (Fithian et al., 2017), where inference is conditioned on those hypothetical datasets which might have occurred which would have led to selection of the same inferential problem; and information splitting methods, such as data splitting, where information in the sample is split between that used for selection and that used for inference: see for example García Rasines and Young (2023); Leiner et al. (2023); Neufeld et al. (2024).
@@
In this work, we study the asymptotic behavior of selection models, which arise as the working inferential models under the conditional and information-splitting approaches to selective inference. These are statistical models in which the data distribution is conditioned on the realization of a certain event. Although selection models have seen renewed attention in recent years, they have long been ubiquitous in statistical inference. Their study dates back at least to Fisher (1934), and many of their foundational mathematical properties were developed by Fraser (1952, 1966). These models have traditionally been used to model situations involving sampling bias, where samples are observed only if they satisfy a certain condition—for example, if they belong to a certain portion of the population (Rao, 1985). Selective inference uses a rather more general formulation, which is described in Section 2.
@@
Within the selective inference literature, much of the methodology for selection models has focused on Gaussian data, typically assuming a known (or consistently estimable) covariance structure. A central reason is that Gaussian models enable a natural treatment of nuisance parameters through standard conditioning arguments; see e.g. Lee et al. (2016). As a result, the theoretical properties of selection models within this setting are well understood, both from the frequentist and the Bayesian positions. The main contribution of this work is the identification of an asymptotic connection between general parametric selection models and Gaussian ones, thereby providing a uniform framework for analyzing the properties of the former and deriving powerful theoretical insights from studying the latter.
@@
More broadly, this paper contributes to the literature on non-standard asymptotics by establishing a limiting behavior which deviates from the standard Gaussian framework. We show that, under mild regularity conditions, a sequence of selection models from a parametric distribution is asymptotically equivalent to a sequence of selection models derived from a Gaussian location model, which, in contrast to standard asymptotics, can be highly asymmetrical if the selection effect is significant (Section 3). The relevant Gaussian models inherit a selection mechanism for the mean parameter that often depends on both the original model and the underlying data-generating parameter in an intricate manner, as illustrated in the examples of Sections 3.4 and 3.5. This showcases the unconventional nature of the analysis, compared to standard asymptotic theory, and underscores the need for careful consideration of the appropriate asymptotic framework for inference.
@@
A key motivation for the study of Local Asymptotic Normality is that analysis of the simpler approximating Gaussian model yields information about asymptotic properties in the original parametric model. In this paper, we focus on the methodological implications for Bayesian inference in selection models. As a natural corollary of the asymptotic approximation, we derive an extension of the Bernstein-von Mises Theorem to selection models (Section 4). We show that in a selection model with a fixed prior on the parameter, the Bayesian posterior is asymptotically equivalent to that obtained in a Gaussian selection model with a uniform prior on the location parameter. Crucially, this is a non-Gaussian limit, with significant practical consequences for selective inference, as the posterior in such a Gaussian selection model is mis-calibrated from a frequentist perspective. In these models, frequentist calibration requires specification of a prior density which can depend strongly on the sample size, as discussed in García Rasines and Young (2022).
2 Framework
The term selection model may refer to slightly different types of statistical models, depending on the context. For the purposes of this work, a selection model is defined by two elements: a base model, understood as the pre-selection model and assumed to contain the true data-generating distribution; and a conditioning random variable, upon which the data is conditioned to obtain the selection model. This variable encodes the selection mechanism, i.e. the rule determining which realizations from the base model are observed, or selected, for some purpose. Operationally, a sample $y^{n}$ is retained only if a certain value of the conditioning variable is observed. This framework is sufficiently general to encompass all applications in the conditional and information-splitting approaches to selective inference, as illustrated in Section 3.
Typically, the base model is well understood from an inferential standpoint, and the objective is to extrapolate its properties to the selection model by analyzing how the selection mechanism interacts with it. In this paper, we will restrict attention to base models consisting of random samples from low-dimensional, regular parametric families. Other classes of base models of central importance in selective inference which are not covered in this work include fixed-design regression, high-dimensional, and non-parametric models.
2.1 Base model
@@
Let $Y^{n}=[Y_{1},\ldots,Y_{n}]$ be a vector of independent and identically distributed (IID) random variables from a distribution $F_{\theta}$ on a measurable space $(\mathcal{Y},\mathcal{A})$, where $\theta$ is a parameter from an open set $\Theta\subseteq\mathbb{R}^{p}$. The model for the full sample, defined over $(\mathcal{Y}^{n},\mathcal{A}^{n})$, is $\mathcal{M}_{n}=\{F_{\theta}^{n}\colon\theta\in\Theta\}$. We regard $\{\mathcal{M}_{n}\colon n\geq 1\}$ as the sequence of base models, from which a sequence of selection models will be obtained. It is well known that, under mild regularity conditions, $\mathcal{M}_{n}$ can be asymptotically approximated by the Gaussian location model $\{N(h,I_{\theta}^{-1})\colon h\in\mathbb{R}^{p}\}$, where $I_{\theta}$ is the Fisher information for a single sample. Specifically, if the model $\{F_{\theta}:\theta\in\Theta\}$ is differentiable in quadratic mean (DQM; van der Vaart, 1998, p. 93), then the sequence $\{\mathcal{M}_{n}:n\geq 1\}$ is Locally Asymptotically Normal (LAN) in the sense of Le Cam (1960). That is, the sequence of local log-likelihood ratios admits the expansion
@@
$\log\prod_{i=1}^{n}\frac{f_{\theta+h_{n}/\sqrt{n}}(Y_{i})}{f_{\theta}(Y_{i})}=h^{\top}I_{\theta}Z_{n}-\tfrac{1}{2}h^{\top}I_{\theta}h+o_{F_{\theta}}(1)\quad\text{as }n\to\infty,$
@@
for every converging sequence $h_{n}\to h$, where $Z_{n}\rightsquigarrow N(0,I_{\theta}^{-1})$ under $F_{\theta}$, and $f_{\theta}(y)$ is the density of $F_{\theta}$ with respect to some fixed measure $\mu$. We use $\rightsquigarrow$ to denote convergence in distribution. In the limit, the local likelihood ratios are distributed as the likelihood ratios of $\{N(h,I_{\theta}^{-1})\colon h\in\mathbb{R}^{p}\}$. This translates into an asymptotic equivalence between both models in likelihood-based inference. Many key asymptotic results can be derived from the LAN property, including the asymptotic distribution of maximum likelihood estimators and the Bernstein–von Mises theorem in Bayesian inference; see van der Vaart (1998) for a detailed account. Throughout this work, we shall assume that $\{F_{\theta}:\theta\in\Theta\}$ is differentiable in quadratic mean. This ensures that the sequence of base models is LAN, which forms the foundation of the asymptotic theory for selection models.
Example 2.1.
We will illustrate the framework through an exponential base model, where $Y_{i}\sim\text{Exp}(\theta)$ and $\Theta=(0,\infty)$. Upon setting $Z_{n}=\theta^{2}\sqrt{n}(1/\theta-\bar{Y}_{n})$ to be the normalized score at $\theta$, where $\bar{Y}_{n}$ is the sample mean, the LAN expansion of $\mathcal{M}_{n}$ specializes to
$\log\prod_{i=1}^{n}\frac{f_{\theta+h_{n}/\sqrt{n}}(Y_{i})}{f_{\theta}(Y_{i})}=\sqrt{n}h\left(\frac{1}{\theta}-\bar{Y}_{n}\right)-\frac{h^{2}}{2\theta^{2}}+o_{F_{\theta}}(1)\quad\text{as }n\to\infty.$
2.2 Selection model
@@
A selection model can be obtained from a base model $\mathcal{M}_{n}$ by conditioning the data on a random quantity $U_{n}$ such that $U_{n}\mid Y^{n}=y^{n}$ is distribution-constant for every $y^{n}$. This implies that $U_{n}$ is either a statistic, such as $u(Y^{n})=\mathbf{1}(\bar{Y}_{n}>0)$, where $\mathbf{1}(A)$ denotes the indicator function of the event $A$, or a randomized function of the data, $u(Y^{n},W_{n})$, say, where $W_{n}$ is user-generated noise not depending on $\theta$. Although this might seem somewhat artificial, randomizing the selection step is standard practice in selective inference. The purpose is to increase the information about $\theta$ contained in the selection model, in a way analogous to data splitting. For a realized value $U_{n}=u_{n}$, the conditional density of $Y^{n}\mid u_{n}$ is
@@
where $f(u_{n}\mid y^{n})$ and $f_{\theta}(u_{n})$ are the conditional and marginal densities of $U_{n}$ relative to some dominating measures, and $f_{\theta}(u_{n})$ is assumed to be positive for every value of $\theta$. We denote the corresponding probability distribution by $F_{\theta}^{u_{n}}$, and define the selection model with base model $\mathcal{M}_{n}$, conditioning variable $U_{n}$, and realized $U_{n}=u_{n}$ as $\mathcal{M}_{n}(u_{n})=\{F_{n}^{u_{n}}\colon\theta\in\Theta\}$. Furthermore, we define a Gaussian selection model as any selection model whose base model is a Gaussian location model: $\mathcal{M}_{n}=\{N(\theta,\Sigma)^{n}\colon\theta\in\mathbb{R}^{p}\}$, where $\Sigma$ is known and positive definite. We will show that Gaussian selection models arise as asymptotic first-order approximations of general selection models under suitable regularity conditions.
Our objective is to understand the asymptotic behavior of $\{\mathcal{M}_{n}(u_{n})\colon n\geq 0\}$ for fixed sequences of $U_{n}$-realizations $\{u_{n}\colon n\geq 1\}$. For this reason, we will occasionally suppress $u_{n}$ from the notation. In particular, we will write
@@
where $p_{n}(y^{n})=f(u_{n}\mid y^{n})$ and $\varphi_{n}(\theta)=f_{\theta}(u_{n})$. This simplifies the expressions, in particular in Section 4 where we discuss Bayesian selective inference, and aligns better with the notation conventionally employed in most selective inference applications. Henceforth, we will refer to $p_{n}(y^{n})$ as the selection function. This encodes all relevant distributional information about the selection mechanism, in the sense that any two statistics and realized values $U_{n1}=u_{n1}$ and $U_{n2}=u_{n2}$ with proportional selection functions $p_{n1}(y^{n})\propto p_{n2}(y^{n})$ induce the same selection model on $Y^{n}$. In many applications, $U_{n}$ is a discrete random variable which indicates whether a certain event of interest $A$ has occurred: $U_{n}=\mathbf{1}(A)$. In such cases, $A$ is called the selection event and, for $u_{n}=1$, $\varphi_{n}(\theta)=P_{\theta}(A)$ is the selection probability. Note, however, that whenever the selection function is bounded, the same selection model can be represented as one induced by a discrete selection variable. Indeed, we may normalize $p_{n}$ to obtain a genuine conditional probability by defining the equivalent selection function
$\tilde{p}_{n}(y^{n})=\frac{p_{n}(y^{n})}{\|p_{n}\|_{\infty}},\quad\|p_{n}\|_{\infty}=\sup\{p_{n}(y^{n})\colon y^{n}\in\mathcal{Y}^{n}\},$
so that $0\leq\tilde{p}_{n}(y^{n})\leq 1$ for all $y^{n}$. We can then introduce an auxiliary binary variable by letting $\tilde{U}_{n}\mid y_{n}\sim\text{Bernoulli}\{\tilde{p}_{n}(y^{n})\}$, so that conditioning on $\tilde{U}_{n}=1$ yields a selective model for $Y^{n}$ that is equivalent to the original one.
@@
Example 2.1 (ctd.).
Consider the exponential base model of Example 2.1 and assume that, for scientific reasons, we are only interested in $\theta$ if it is large. That is, we would like to perform inference on $\theta$ only if the sample information indicates a large enough rate $\theta$, according to some measure of calibration. Let $U_{n}=\mathbf{1}(\sqrt{n}\bar{Y}_{n}+W<t)$ for some threshold $t$, where $W\sim N(0,1)$ independently of $Y^{n}$, and assume a sample is selected for further analysis only if $U_{n}=1$, which may be regarded as rejecting the null hypothesis in a randomized one-sided test for $\theta$. The conditional approach to selective inference dictates that inference following observation of $U_{n}=1$ ought to be based on the corresponding selection model, with respective selection function and probability
$p_{n}(y^{n})$ $=$ $\Phi\left(t-\sqrt{n}\bar{y}_{n}\right);$ (3)
$\varphi_{n}(\theta)$ $=$ $E_{\theta}\left[\Phi\left(t-\sqrt{n}\bar{Y}_{n}\right)\right];$ (4)
where $\Phi$ is the $N(0,1)$ distribution function.
Suppose we wish to test $H_{0}\colon\theta\leq\theta_{0}$ against the alternative $H_{1}\colon\theta>\theta_{0}$ for large $n$. An exact test under the selection model requires comparison of the observed $\bar{y}_{n}$ with the sampling distribution of $\bar{Y}_{n}$ when $\theta=\theta_{0}$, given selection. Analytically, this requires consideration of the convolution of a Gamma and a Gaussian distribution, which in this case is analytically complicated, though easily simulated. In the following section we will show local asymptotic equivalence, at a given $\theta$, between $\mathcal{M}_{n}(1)$ and a Gaussian selection model. The test of $H_{0}$ can therefore be approximated by referring the observed value of $z_{n}(\theta_{0})=\theta_{0}^{2}\sqrt{n}(1/\theta_{0}-\bar{y}_{n})$ to the asymptotic density
$f_{\theta_{0}}(z)=\frac{1}{\theta_{0}}\cdot\frac{\phi(\frac{z}{\theta_{0}})\Phi\bigl(t-\sqrt{n}/\theta_{0}+\frac{z}{\theta_{0}^{2}}\bigr)}{\Phi\big\{(t-\sqrt{n}/\theta_{0})/\sqrt{1+1/\theta_{0}^{2}}\bigr\}},\quad z\in\mathbb{R},$ (5)
where $\phi$ is the standard Gaussian density. Evidence against $H_{0}$ is provided by small values of $z_{n}(\theta_{0})$. The appropriate approximate $p$-value is now easily evaluated by a simple numerical integration of the density (5).
Such use of the asymptotic Gaussian selection model to examine repeated sampling properties in the initial exponential selection model is illustrated in Figure 1 for two sample sizes, $n=50,200$. For each sample size, we consider, over a series of 10,000 replications, the distribution in the two models of the $p$-value for testing $H_{0}\colon\theta\leq 2$ against the alternative $H_{1}\colon\theta>2$, under two scenarios, when $H_{0}$ is true, and when the alternative $H_{1}$ is true, with $\theta=2.15$. In each case we calculate the $p$-value under the true selection model by simulating the null sampling distribution of $\bar{Y}_{n}$ given the selection condition $\sqrt{n}\bar{Y}_{n}+W<t$ from a series of 100,000 samples. The asymptotic $p$-values are obtained by numerical integration of the density (5). The selection thresholds are set as $t=2.1$ and $t=5.6$ respectively for the two sample sizes $n=50,200$: these thresholds correspond to low selection probability $\approx 10\%$ under the null. Figure 1 shows that even for $n=50$ the two $p$-value distributions are close under both the null and alternative. The distributions are quite indistinguishable for the larger sample size: repeated sampling behaviour under the original randomized selection model is closely matched by that seen in the simpler asymptotic selection model. Inferential properties under the true selection model are accurately predicted from those of the asymptotic approximating Gaussian selection model.
Figure 1: Repeated sampling distribution of $p$-values under the true randomized exponential selection model (blue) and the asymptotic Gaussian selection model (black), $n=50,200$, under null and alternative.
In the Bayesian setting, we will show that the repeated-sampling behavior of posterior distribution for $\theta$ in the selection model can be derived from the asymptotically equivalent Gaussian model. This reveals a fundamental divergence between the frequentist and Bayesian approaches in selection models that contrasts with classical theory for IID models.